OpenStudy (anonymous):

solve equation: 51=7(-1+2v)+2

OpenStudy (anonymous):

what are you having a problem with?

OpenStudy (anonymous):

do you know what the first step is?

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

im freaking out

OpenStudy (anonymous):

subtract 2 from 51 you would then get 49=7(-1+2v) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know what to do next?

OpenStudy (anonymous):

hold on let me think

OpenStudy (anonymous):

do you distribute?

OpenStudy (anonymous):

no you do division... so what you would do now is divide 49 by 7... what is that equal to?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

good so what do you have now?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i have 7=-1+2v

OpenStudy (anonymous):

i believe they should distribute which would give them 49=-7+14v. Then add 7 and get 56=14v. Then divide and there is your answer.

OpenStudy (anonymous):

you still get the same answer

OpenStudy (anonymous):

ok, i get what your saying. EIther way works.